The distribution of height of adult male of a particular race is normally distributed with




mean 165 cm and variance 36 cm. find the probability of an adult male to have height




between 150 cm and 180 cm. [given that 0 ≤ Z ≤ 2.5 = 0.4938]

1
Expert's answer
2021-12-01T18:03:52-0500

"\mu = 165 \\

\sigma^2 = 36 \\

\sigma= 6 \\

P(150<X<180) = P(X<180) -P(X<150) \\

= P(Z< \frac{180-165}{6}) -P(Z< \frac{150-165}{6}) \\

= P(Z<2.5) -P(Z< -2.5) \\

= 0.9937 -0.0062 \\

= 0.9875"